OpenStudy (anonymous):

Type the second number for each pair such that square root of six is the geometric mean for the pair. What is the second number in the pair if the first number is 2?

OpenStudy (anonymous):

The geometric mean of two numbers a and b is \[\sqrt{ab}\]. So solve \[\sqrt{2a}=6\].

OpenStudy (anonymous):

woops i meant solve \[\sqrt{2a}=\sqrt{6}\]